. In order to fabricate a uniform wire of resistance 2.7 ?, out of m=1 g...

.

In order to fabricate a uniform wire of resistance 2.7 ?, out of m=1 g copper which has a density of 8.92
Answer #1

density = mass/volume

volume (AL) = mass/density = 10^-3 /8.92 *10^3   = 11.2*10^-8


also Resisatnce R = Rho L/A

L/A = R/Rho = 2.7/1.7*10^-8 = 1.588*10^8


multiply L/A and AL

L^2 = 17.7856

L = 4.212 m



divide LA/L/A =A^2 = 11.2*10^-8/1.588*10^8

A^2 = 7.05 *10^-16

A = 2.65 *10^-8

pi r^2 = 2.65*10^-8

r^2 = 0.845 *10^-8

r = 0.919 *10^-4 m

diameter d = 2r = 1.83 *10^-4 m
Answer #2

a)
L = ?[m?/R?]
where m = mass, kg; ? = resistance in ohms; R = resistivity of copper (1.7E-8 ??m); ? = copper density in kg/m^3

b) D = 2?[m/?L?] (D in meters)